### v1

# Understanding Before Recommendation: Semantic Aspect-Aware Review Exploitation via Large Language Models

Abstract.
Recommendation systems harness user-item interactions like clicks and reviews to learn their representations. Previous studies improve recommendation accuracy and interpretability by modeling user preferences across various aspects and intents. However, the aspects and intents are inferred directly from user reviews or behavior patterns, suffering from the data noise and the data sparsity problem. Furthermore, it is difficult to understand the reasons behind recommendations due to the challenges of interpreting implicit aspects and intents. Inspired by the deep semantic understanding offered by large language models (LLMs), we introduce a chain-based prompting approach to uncover semantic aspect-aware interactions, which provide clearer insights into user behaviors at a fine-grained semantic level. To incorporate the abundant interactions of various aspects, we propose the simple yet effective Semantic Aspect-based Graph Convolution Network (short for SAGCN). By performing graph convolutions on multiple semantic aspect graphs, SAGCN efficiently combines embeddings across multiple semantic aspects for final user and item representations. The effectiveness of the SAGCN was evaluated on three publicly available datasets through extensive experiments, which revealed that it outperforms all other competitors. Furthermore, interpretability analysis experiments were conducted to demonstrate the interpretability of incorporating semantic aspects into the model.
Recommendation, Semantic Aspect, Large Language Models, Collaborative Filtering
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Information systems Personalizationccs: Information systems Recommender systemsccs: Information systems Collaborative filtering
1. Introduction
Recommendation systems have become essential techniques for various online platforms. It not only assists users in navigating through overwhelming volumes of information but also serves as a powerful tool for service providers to increase customer retention and revenue. Within this field, Collaborative Filtering (CF)-based models have achieved significant advancements in learning the user and item representations by effectively exploiting the historical user-item interactions (Bell and Koren, 2007; He et al., 2017; Wu et al., 2016; Xue et al., 2017). For instance, Matrix Factorization (MF) decomposes the user-item interaction matrices into lower-dimensional latent vectors and predicts the user preference with the inner product (Bell and Koren, 2007). By replacing the inner product with nonlinear neural networks, NeuMF (He et al., 2017) can model the complex interaction behaviors, yielding better recommendation performance. More recently, Graph Convolution Networks (GCNs)-based models (van den Berg et al., 2018; Wang et al., 2019b, a; He et al., 2020; Wang et al., 2020a; Liu et al., 2021; Fan et al., 2022) offer a cutting-edge approach in recommendation systems by leveraging the high-order connectivities based on the topological structure of the interaction data. For example, NGCF (Wang et al., 2019b) incorporates high-order connectivity of user-item interactions in the user-item bipartite graph for representation learning in recommendation.
[…]
Semantic aspects often denote specific features or attributes of a product or service in which users might be interested in the context of recommendation systems and sentiment analysis. Extracting semantic aspects from reviews could enhance our comprehension of the complex relations between each user and item at a finer level. Moreover, semantic aspects, like meaningful attributes or concepts of an item, provide a deeper understanding of its inherent features, thereby enhancing the interpretability of recommendations. As shown in Fig. 1, the review might mention various semantic aspects such as functionality, durability, and ease of use. Each of these aspects carries a specific meaning related to user experience. By analyzing and summarizing the raw review into structured semantic aspects, we can better understand the underlying rationales behind user behaviors. Focusing on reviews that highlight specific item aspects enables us to discern semantic aspect-aware interactions. For instance, a review might emphasize aspects like functionality and durability but omit ease of use. Such patterns indicate distinct interactions between the user and item concerning functionality and durability, but not for ease of use. Leveraging these semantic-aware insights leads to richer and more accurate representations of both users and items.
Figure 1. An example of extracting semantic aspect-aware reviews (e.g., functionality, durability, and ease of use) from user review via LLM. Based on the extracted reviews, the semantic aspect-aware interactions are discerned.
Large Language Models (LLMs) have emerged as powerful tools in natural language processing and have recently gained significant attention in various fields (Da’u and Salim, 2022; Sileo et al., 2022). Their contextual understanding allows for accurately distilling various semantic aspects based on reviews without being affected by noise present in reviews. Moreover, based on the extracted aspects, we can form aspect-specific interactions, capturing users’ preferences from various semantic aspects. Motivated by the above consideration, we devise a chain-based prompting approach, which consists of the semantic aspects extraction prompt and semantic aspects-aware review extraction prompt. The former prompt extracts semantic aspects from raw reviews, while the latter one identifies reviews aligned with these distinct semantic aspects. To demonstrate the potential of structuring raw reviews into semantic aspects via LLMs in recommendations, we introduce a Semantic Aspect-based Graph Convolution Network (SAGCN) model. This model performs graph convolution operations on aspect-aware interaction graphs, each of which is constructed based on the user-item interactions of a specific aspect derived from the aspect-aware reviews. The final user and item representations are obtained by combining the embeddings learned from different aspect-specific graphs. The merits of our model are twofold: (1) embeddings are enriched by learning from various aspects; (2) interpretability is improved by incorporating semantic aspects. Extensive experiments are conducted on three benchmark datasets to evaluate the SAGCN. Results show that SAGCN can gain better user and item representations than SOTA GCN-based recommendation models. The interpretability analysis experiments demonstrate the interpretability of incorporating semantic aspects into our model.
In summary, the main contributions of this work are as follows:
  • We emphasize the significance of utilizing semantic aspect-aware interaction in recommendations. Inspired by this, we propose a chain-based prompting approach to extract the semantic aspect-aware user-item reviews, which are used to discern aspect-aware interactions.
  • We introduce an SAGCN model to leverage the extracted semantic aspects via LLM, demonstrate the effectiveness of the extracted semantic aspects by our method on enhancing the recommendation accuracy and interpretability.
  • We conducted empirical studies on three benchmark datasets to evaluate the performance of SAGCN. Our results demonstrate that SAGCN outperforms state-of-the-art GCN-based recommendation models in terms of accuracy and interpretability.
2. related work
2.1. Collaborative Filtering
Collaborative Filtering (CF) techniques (Koren et al., 2009; Rendle et al., 2009; Wang et al., 2019a; He et al., 2017; Wang et al., 2019b; He et al., 2020; Liu et al., 2019) play an crucial role in recommendation systems. The core idea is to learn users’ and items’ representations ($aka.$ embedding vectors) by exploiting the collaborative information based on user-item interactions. With the learned user and item embedding vectors, an interaction function is adopted to predict a user’s preference for an item. For example, Matrix Factorization (MF) (Koren et al., 2009) learns the user and item embeddings by minimizing the error of re-constructing the user-item interaction matrix. Later, due to its powerful capability, the deep learning technique is widely adopted in recommendations to learn better user and item embeddings (Xue et al., 2017) or model more complicated interactions between users and items (He et al., 2017; Li et al., 2019). For example, NeuMF (He et al., 2017) leverages a nonlinear neural network to model the complex interactions between users and items. Owing to the powerful ability to learn representations based on the graph structure, graph convolution networks are employed to exploit the high-order user-item interactions for improving recommendation performance (van den Berg et al., 2018; Wang et al., 2019b). A typical method is NGCF (Wang et al., 2019b), which aggregates embeddings from high-order neighbors by performing graph convolution on the user-item interaction graph.
[…]
2.2. GCN-based Recommendation Models
In recent years, Graph Convolutional Networks (GCNs) have garnered significant attention in the realm of recommendation systems, owing to their robust capabilities in learning node representations from non-Euclidean structures (van den Berg et al., 2018; Wang et al., 2019b; Ying et al., 2018; Wang et al., 2019a, 2020a; He et al., 2020; Liu et al., 2021; Lin et al., 2022; Liu et al., 2022b; Cheng et al., 2023). This also led to the rapid advancement of GCN-based recommendation models. For example, GC-MC (van den Berg et al., 2018) employs one convolution layer to exploit the direct connections between users and items. NGCF (Wang et al., 2019b) propagates embeddings on the user-item interaction graph to exploit the high-order proximity. Inspired by the study of simplifying GCN (Wu et al., 2019), researchers also introspect the complex design in GCN-based recommendation models. For example, He et al. (He et al., 2020) observed that the two common designs feature transformation and nonlinear activation have no positive effects on the final performance. Thus, they proposed LightGCN, which removes these two modules and significantly improves recommendation performance.
To achieve robust node representation, fine-grained interactions are exploited to refine the graph structure. Wang et al. (Wang et al., 2020a) presented a GCN-based model, which yields disentangled representations by modeling a distribution over intents for each user-item interaction; IMP-GCN (Liu et al., 2021) learns user and item embeddings by performing high-order graph convolution inside subgraphs constructed by the user nodes with common interests and associated items. By incorporating side information, GRCN (Wei et al., 2020) adjusts the structure of the interaction graph adaptively based on the status of model training for discovering and pruning potential false-positive edges; and LATTICE (Zhang et al., 2021) mines the latent structure between items by learning an item-item graph. Graph convolutional operations are performed on both item-item and user-item interaction graphs to learn user and item representations.
2.3. LLMs in Recommendation
Large Language Models (LLMs) have emerged as powerful tools in natural language processing and have recently gained increasing attention in recommendation systems. The LLM-based models are particularly proficient at capturing contextual information and comprehending user queries, item descriptions, and other textual data in a highly effective manner (Geng et al., 2022; Da’u and Salim, 2022; Sileo et al., 2022; Wu et al., 2023). By understanding the context, LLM-based models can improve the recommendation performance.
There are two main categories for incorporating LLMs into recommendation systems. One is to adapt LLMs as recommendation models. Most generative model-based work translates recommendation tasks as natural language tasks and then applies techniques such as in-context learning, prompt tuning, and instruction tuning to adapt LLMs to generate the recommendation results (Ji et al., 2023; Zhang et al., 2023b, a; Kang et al., 2023; Li et al., 2023). Moreover, with the impressive capabilities demonstrated by ChatGPT, this type of work has received increased attention recently. Ji et al. (Ji et al., 2023) proposed a generative recommendation method that uses LLMs to generate recommended items. The other is to extract high-quality representations of textual features or leverage the extensive external knowledge encoded within them (Liu et al., 2023). Most existing works align the representations of pre-trained models like BERT (Devlin et al., 2019) with the domain-specific data through fine-tuning. In addition to taking LLMs as recommendation systems, some studies also utilize LLMs to construct model features. For example, GENRE (Liu et al., 2023) introduced three prompts to conduct three feature enhancement sub-tasks for news recommendation. The prompts involved refining news titles using ChatGPT, extracting profile keywords from user reading history, and generating synthetic news to enhance user interactions based on their history. By incorporating these LLM-constructed features, traditional news recommendation models can be greatly improved.
3. Methodology
3.1. Preliminary
3.1.1. Problem Setting and Notations Definition.
Before describing our model, we would like to introduce the problem setting first. A dataset with an interaction matrix $\bm{R}^{N\times M}$ of a set of users $\mathcal{U}$ and a set of items $\mathcal{I}$ is involved, where $N$ and $M$ are the numbers of users and items, respectively. The matrix $\bm{R}$ records the interaction history between users and items. A nonzero entry $r_{ui}\in\bm{R}$ indicates that user $u\in\mathcal{U}$ has interacted with item ${i}\in\mathcal{I}$ before; otherwise, the entry is zero. For each user-item pair $(u,i)$, a review $v$ is associated with them. The review contains the user’s sentiment on different aspects of an item. We employ LLM to extract a semantic aspect set $\mathcal{A}$, which includes all the aspects mentioned in the reviews. Specifically, $\mathcal{A}_{v}=\{a_{v}^{1},\cdots,a_{v}^{N_{v}}\}$ is the set of semantic aspects extracted from review $v$, where $N_{v}$ is the number of extracted aspects. In this work, our goal is to explore the potential of the extracted semantic aspects from raw reviews by LLMs to enhance recommendation accuracy and interpretability.
To achieve the goal, we (1) design a chain-based prompting approach, which can effectively extract semantic aspect-aware reviews automatically; (2) introduce a semantic aspect-based GCN model for recommendation. We use $\{\bm{R^{1}},\bm{R^{2}},\cdots,\bm{R^{A}}\}$ to denote the semantic aspect-based interaction matrices. $A$ is the number of aspects. $\bm{R^{a}}$ is the interaction matrix of the $a$-th aspect.
Figure 2. Chain-based Prompting Approach.
3.1.2. Semantic Aspect-based Graph Construction.
The semantic aspect-based graph construction is built on the results of aspect-aware review extraction, which is achieved by two steps with a designed chain-based prompting approach. We extract semantic aspects using the semantic aspects extraction prompt and then extract aspect-aware reviews based on those aspects using the semantic aspects-aware review extraction prompt.
Semantic aspect-aware interaction: if a user $u$’s review contains comments of a specific on a certain aspect of item $i$, then we regard there is an interaction between u and i on this aspect. By identifying the aspects of user reviews on each item, we can form semantic aspect-aware interaction graphs to exploit the semantic aspects in user and item representation learning.
In previous studies (Wang et al., 2019b; He et al., 2020), the user-item interaction graph is directly built based on the users’ historical interactions. Specifically, let $\mathcal{G}=(\mathcal{W},\mathcal{E})$ be an undirected graph, where $\mathcal{W}$ denotes the set of nodes and $\mathcal{E}$ is the set of edges. Specifically, $\mathcal{W}$ consists of two types of nodes: users $u_{i}\in\mathcal{U}$ with $i\in\{1,\cdots,N_{u}\}$, items $v_{j}\in\mathcal{V}$ with $j\in\{1,\cdots,N_{v}\}$, and $\mathcal{U}\cup\mathcal{V}=\mathcal{W}$. By extracting the semantic aspect-aware interactions from reviews using LLM, we separately construct the semantic aspect-specific graphs based on the user-item interactions on various aspects. In this work, we defined $\mathcal{G}_{a}=(\mathcal{W},\mathcal{E}_{a})$ as the semantic aspect-based graph for aspect $a$, where $\mathcal{E}_{a}$ is the set of edges.
In the next, we detail our method for semantic aspect-aware review extraction.
3.2. Semantic Aspect-Aware Review Extraction
In this section, we conduct sentiment analysis on user reviews using LLM to discern user-item interactions on various semantic aspects. As illustrated in Fig 2, we introduce a chain-based prompting approach by combining the semantic aspects and semantic aspect-aware review extraction.
3.2.1. Semantic Aspects Extraction
The semantic aspects vary depending on the recommendation scenario. For instance, when buying dresses, style is crucial, but for office products, ease of use is the main concern. Empirically, it is challenging to extract the high-quality unseen aspects and reviews simultaneously. Therefore, we first extract semantic aspects from all reviews of each dataset. The semantic aspects extraction prompt is designed as follows:
  • Prompt 1: A person bought a product and commented that ¡review¿. Tell me from which perspectives the customer gave this review, e.g., quality, comfort, etc. Answer point by point.
As shown in Fig. 1, LLM generates a semantic aspect set and its corresponding analysis for each aspect. Note that, we provide little information about the semantic aspect in our prompt, leading to low-quality analysis. By iteratively analyzing each review using LLM, we can obtain all semantic aspects within this recommendation domain by aggregating all the semantic aspect sets of all reviews.
3.2.2. Semantic Aspect-Aware Review Extraction
Based on the extracting semantic aspects, we further employ LLM to generate high-quality semantic aspect-aware reviews. By focusing on specific aspects, high-quality semantic aspect-aware reviews can be obtained. For instance, on the Baby Dataset, we get a semantic aspect set containing Quality, Comfort, Durability, Design, Functionality, Ease of use, Price, and Size. To accurately identify reviews to different aspects, we adapt the prompt to the semantic aspects of the recommendation scenario:
  • Prompt 2: A person bought a product and commented that ¡review¿. Tell me from which perspectives the customer gave this review, e.g., Quality, Comfort, Durability, Design, Functionality, Ease of Use, Price, and Size. Answer point by point.
The LLM identifies the semantic aspect-aware review for the aspects we input. Then, we identify the semantic aspect-aware interactions between the user and the item based on the generated semantic-aware reviews. For example, if the LLM generates the output: “The customer mentioned that the bins were not durable and that the fabric disintegrated when burned with a candle.”, it indicates an interaction on Functionality. On the other hand, if the LLM generates the output: “The customer did not mention anything about the ease of use of the bins.”, which means that there is no interaction on this aspect.
3.3. Semantic Aspect-Based Graph Convolution Network
Figure 3. Overview of our SAGCN model.
3.3.1. Overview
Based on the extracted semantic-aware user-item interactions for semantic aspects, we seek to model users’ preferences on various aspects with GCN-based techniques. The structure of our SAGCN is illustrated in Figure 1, and it is composed of three key components: 1) embedding initialization, which initialized node (users and items) embeddings for the subsequent learning in each aspect-specific interaction graph; 2) embedding propagation, in which LightGCN (He et al., 2020) is adopted to learn the user and item embeddings by performing graph convolution on each semantic aspect-based graph; 3) embedding generation, which aggregates the embeddings learned from each aspect-specific graph and combines the embeddings of different aspects.
3.3.2. Embedding Initialization
In our model, the vector of users and items is composed of embeddings from A aspects. Following existing GCN-based recommendation models (Wang et al., 2019b; He et al., 2020; Liu et al., 2021), we initialize the embedding vectors for each aspect. As an example, the ID embeddings of user $u\in\mathcal{U}$ for different aspects are initialized as:
(1) $\bm{e_{u}}^{(0)}=\{\bm{e_{u}^{(1,0)},e_{u}^{(2,0)},\cdots,e_{u}^{(A,0)}}\},$
where $\bm{e_{u}^{(0)}}\in\mathds{R}^{d\times A}$ is ID embedding to capture intrinsic characteristics of user $u$; $\bm{e_{u}^{(a,0)}}\in\mathds{R}^{{d}}$ is $u$’s representation in aspect $a$. Analogously, $\bm{e_{i}}^{(0)}=\{\bm{e_{i}^{(1,0)},e_{i}^{(2,0)},\cdots,e_{i}^{(A,0)}}\}$ is defined as the ID embeddings of item $i\in\mathcal{I}$. The embeddings of the user ID and item ID for each aspect are randomly initialized in our model.
3.3.3. Embedding Propagation
To exploit the high-order connectives on each aspect to learn user and item representations, we propagate embeddings on multiple semantic aspect-based graphs. Take the semantic aspect-specific graph $\mathcal{G}_{a}$ as an example, let $\bm{e_{u}^{(a,k)}}$ and $\bm{e_{i}^{(a,k)}}$ denote the embedding of user $u$ and item $i$ after $k$ layers graph convolution, where $k\geq 0$, the embedding propagation is formulated as:
(2) $\bm{e_{u}^{(a,k+1)}}=\sum_{i\in\mathcal{N}_{u}^{a}}\frac{1}{\sqrt{|\mathcal{N}_{u}^{a}|}\sqrt{|\mathcal{N}_{i}^{a}|}}\bm{e_{i}^{(a,k)}},$
$\bm{e_{i}^{(a,k+1)}}=\sum_{u\in\mathcal{N}_{i}^{a}}\frac{1}{\sqrt{|\mathcal{N}_{i}^{a}|}\sqrt{|\mathcal{N}_{u}^{a}|}}\bm{e_{u}^{(a,k)}},$
where $\bm{e_{i}^{(a,k+1)}}$ and $\bm{e_{u}^{(a,k+1)}}$ can be considered as embeddings propagated from the $(k+1)$-th order neighboring nodes. $\frac{1}{\sqrt{|\mathcal{N}_{u}^{a}|}\sqrt{|\mathcal{N}_{i}^{a}|}}$ is symmetric normalization terms; $\mathcal{N}_{u}^{a}$ denotes the set of items that interact with user $u$, and $\mathcal{N}_{i}^{a}$ denotes the set of users that interact with item $i$.
3.3.4. Embedding Combination
After the $K$-layers graph convolution, the embeddings of user $u$ and item $i$ of aspect $a$ are aggregated from all layers:
(3) $\bm{e_{u}^{(a)}}=\sum_{k=0}^{K}\bm{e_{u}^{(a,k)}},\bm{e_{i}^{(a)}}=\sum_{k=0}^{K}\bm{e_{i}^{(a,k)}}.$
The final representations of user $u$ and item $i$ are a concatenation of their embeddings learning in multiple semantic aspect-based graphs:
(4) $\bm{e_{u}}=\bm{e_{u}^{(0)}}||\cdots||\bm{e_{u}^{(a)}}||\cdots||\bm{e_{u}^{(A)}},\bm{e_{i}}=\bm{e_{i}^{(0)}}||\cdots||\bm{e_{i}^{(a)}}||\cdots||\bm{e_{i}^{(A)}},$
where $\bm{e_{u}^{(a)}}$ and $\bm{e_{i}^{(a)}}$ denote the embeddings of users and items by performing graph convolution on graph $\mathcal{G}_{a}$, respectively.
3.3.5. Model Training
After learning the representation for both users and items, the preference of a given user $u$ for a target item $i$ is estimated as follows:
(5) $\hat{r}_{ui}=\bm{e_{u}}^{T}\bm{e_{i}}.$
The top-$n$ recommendation aims to recommend a set of $n$ top-ranked items matching the target user’s preference. Similar to previous works (Liu et al., 2021), we adopt the pairwise learning method for optimization. To train the proposed model, we construct a triplet of $\{u,i^{+},i^{-}\}$, with a positive item (an observed interaction between $u$ and $i^{+}$) and a negative item (an unobserved interaction between $u$ and $i^{-}$ [^1] Note that pairwise learning assumes that a positive item (i.e., $i^{+}$) should rank higher than a negative item (i.e., $i^{-}$).. The objective function is formulated as follows:
(6) $\mathop{\arg\min}\sum_{(\mathbf{u},\mathbf{i}^{+},\mathbf{i}^{-})\in{\mathcal{O}}}-\ln\phi(\hat{r}_{ui^{+}}-\hat{r}_{ui^{-}})+\lambda\left\|\Theta\right\|^{2}_{2}$
where $\mathcal{O}=\{(u,i^{+},i^{-})|(u,i^{+})\in\mathcal{R^{+}},(u,i^{-})\in\mathcal{R^{-}}\}$ denotes the training set; $\mathcal{R^{+}}$ indicates the observed interactions between user $u$ and $i^{+}$ in the training dataset, and $\mathcal{R^{-}}$ is the sampled unobserved interaction set. $\lambda$ and $\Theta$ represent the regularization weight and the parameters of the model, respectively. The $L_{2}$ regularization is used to prevent overfitting. Besides, we adopt the mini-batch Adam (Kingma and Ba, 2015) to optimize the model and update its parameters.
4. Experiments
4.1. Experimental setup
4.1.1. Data Description.
Table 1. Statistics of the Experimental Datasets.
Dataset #user #item #interactions sparsity
Office 4,905 2,420 53,258 99.55%
Baby 19,445 7,050 160,792 99.88%
Clothing 39,387 23,033 278,677 99.97%
Table 2. Performance of our method and the competitors over three datasets.
Datasets Office Baby Clothing
k=10 k=20 k=10 k=20 k=10 k=20
Metrics Recall NDCG Recall NDCG Recall NDCG Recall NDCG Recall NDCG Recall NDCG
NeuMF 5.14 3.89 8.12 5.21 3.11 2.11 4.85 2.69 0.94 0.54 1.50 0.71
GCMC 6.72 5.27 10.27 6.79 4.55 2.99 7.24 3.89 3.17 1.86 4.86 2.35
LightGCN 9.87 6.04 14.47 7.43 5.94 3.30 9.25 4.20 4.45 2.43 6.44 2.95
DGCF 9.95 6.25 14.37 7.52 5.90 3.27 9.20 4.14 4.67 2.68 6.91 3.32
IMP-GCN 10.11 6.36 14.47 7.71 6.24 3.49 9.56 4.38 4.80 2.76 7.11 3.40
NCL 10.07 6.30 14.40 7.65 6.15 3.42 9.45 4.30 4.76 2.74 7.10 3.37
MMGCN 5.74 3.42 9.39 4.54 3.95 2.17 6.46 2.85 2.42 1.29 3.76 1.64
GRCN 10.38 6.34 15.33 7.81 5.57 3.03 8.49 3.83 4.47 2.35 6.70 2.94
LATTICE 10.00 6.09 14.99 7.57 6.06 3.40 9.29 4.27 5.03 2.79 7.28 3.37
BM3 9.80 6.09 14.02 7.36 6.44 3.65 9.52 4.48 5.28 2.94 7.75 3.58
SAGCN 11.71* 7.34* 16.71* 8.84* 7.35* 4.23* 10.56* 5.09* 6.07* 3.58* 8.44* 4.20*
Imporv. 12.81% 15.77% 9.00% 13.19% 14.13% 15.89% 10.46% 13.61% 14.96% 21.77% 8.90% 17.32%
  • The symbol * denotes that the improvement is significant with $p-value<0.05$ based on a two-tailed paired t-test. Notice that the values are reported by percentage with ’%’ omitted.
Amazon Review Datasets  [^2] http://jmcauley.ucsd.edu/data/amazon stems from the extensive product review system of Amazon (Liu et al., 2019). It includes user-product interactions, reviews, and metadata. In our experiments, only the interaction information is utilized. Three product categories from the 5-core version of this dataset are selected for evaluation: Office, Baby, and Clothing.
For each dataset, we randomly split the interaction of each user into training and testing sets with a ratio of 80:20. From the training set, we randomly select 10% of interactions to form the validation set for tuning hyper-parameters. The pairwise learning strategy is adopted for parameter optimization. Specifically, the observed user-item interactions in the training set were treated as positive instances. For each positive instance, we pair it with a random sampling negative instance that the user has not interacted with yet.
4.1.2. Compared Baselines.
To demonstrate the effectiveness of ClusterGCF, we compared it with several baseline models.
  • NeuMF (He et al., 2017): It is a state-of-the-art neural collaborative filtering method that introduces multiple hidden layers to capture the non-linear interactions between their features.
  • GCMC (van den Berg et al., 2018): This method leverages the GCN technique (Kipf and Welling, 2017) on the user-item bipartite graph and employs one convolutional layer to exploit direct connections between users and items.
  • LightGCN (He et al., 2020): This model can be considered a simplified version of NGCF (Wang et al., 2019b), where certain components like feature transformation and nonlinear activation module are removed.
  • IMP-GCN (Liu et al., 2021): This method performs high-order graph convolution within sub-graphs, each of which is constructed based on users with shared interests and associated items, thus avoiding the negative information passed from high-order neighbors.
  • NCL (Lin et al., 2022): This is a model-agnostic contrastive learning framework. It seeks to capture the correlation between a node and its prototype to improve the neural graph collaborative filtering.
  • LATTICE (Fan et al., 2022): This model uses side information to construct an item-item graph and perform graph convolution operations to aggregate information from high-order neighbors.
  • MMGCN (Wei et al., 2019): The model learns user preferences and item characteristics by propagating modality embeddings on the user-item graph.
  • GRCN (Wei et al., 2020): This GCN-based multimedia recommendation model adjusts the structure of the interaction graph adaptively based on the status of model training for discovering and pruning potential false-positive edges.
  • BM3 (Zhou et al., 2023): This self-supervised framework uses a latent representation dropout mechanism to generate the target view of a user or an item. BM3 is trained without negative samples by jointly optimizing three multi-modal objectives.
For fair comparisons, all methods are optimized using the same pairwise learning strategy. We put great efforts into tuning these methods based on the validation dataset and reported their best performance.
4.1.3. Evaluation Metrics
Two widely-used evaluation metrics are adopted for top-$K$ recommendation: Recall and Normalized Discounted Cumulative Gain (NDCG) (He et al., 2015). Recommendation accuracy is calculated for each metric based on the top 10 and 20 results. Note that the reported results are the average values across all testing users.
4.1.4. Parameter Settings
The PyTorch framework  [^3] https://pytorch.org. is adopted to implement the proposed model. Vicuna-13B  [^4] https://huggingface.co/lmsys/vicuna-13b-v1.5 is used to analyze the reviews. In our experiments, all hyperparameters are carefully tuned. The user and item embedding size of each aspect is set to 64 across all datasets. The mini-batch size is fixed to 1024. The learning rate for the optimizer is searched from $\{1e^{-5},1e^{-4},\cdots,1e^{1}\}$, and the model weight decay is searched in the range $\{1e^{-5},1e^{-4},\cdots,1e^{-1}\}$. Besides, the early stopping strategy is adopted. Specifically, the training process will stop if recall@10 does not increase for 20 successive.
Table 3. The extracted semantic-aware aspects and the number of their associated interactions over three datasets.
Datasets Semantic-aware Aspects
Office Quality Functionality Ease of Use Convenience Comfort Durability Design Price
43,850 43,269 42,347 41,238 40,795 37,564 23,973 23,661
Baby Quality Functionality Comfort Ease of Use Design Durability Size Price
133,337 132,827 127,938 125,014 119,013 116,212 86,125 59,859
Clothing Quality Comfort Appearance Style Fit Design Size Price
230,162 210,254 205,532 188,378 186,132 181,222 170,869 106,416
4.2. Performance Comparison
In this section, we compared the performance of our proposed method with all the adopted competitors. The experimental results (in terms of Recall@10, Recall@20, NDCG@10, and NDCG@20) of our method and all compared methods over three test datasets are reported in Table 2. Note that the best and second-best results are highlighted in bold. From the results, we have the following observations.
We first focus on the performance of methods in the first block, which only uses the user-item interaction information. NeuMF is a deep learning-based method, and it shows relatively poor performance. GCMC performs better than NeuMF, highlighting the benefits of leveraging graph structure. However, GCMC underperforms LightGCN because it only utilizes first-order neighbor nodes. This showcases the strength of GCN and the importance of incorporating high-order information into representation learning. Moreover, DGCF performs better than LightGCN by leveraging the disentangled representation technique to model users’ diverse intents. Both IMP-GCN and NCL consistently outperform all the above GCN-based methods over all the datasets. This is mainly due to their ability to filter out noisy information from neighboring nodes effectively. In particular, IMP-GCN’s performance superiority over LightGCN and DGCF demonstrates the importance of distinguishing nodes in high-order neighbors in the graph convolution operation.
The second block is the GCN-based methods utilizing side information. MMGCN is the first method combining side information with GCN, but it was found to be less effective than GRCN. This is attributed to the GRCN adaptively adjusting the false-positive edges of the interaction graph to discharge the ability of GCN. LATTICE outperforms GRCN on Baby and Clothing by constructing item-item graphs based on item similarity in the side information. BM3 consistently performs better than all general CF baselines over Baby and Clothing. It should be credited to the joint optimization of multiple objectives to learn the representations of users and items.
SAGCN consistently yields the best performance across all datasets. In particular, the improvements of SAGCN over the strongest competitors $w.r.t$ NDCG@10 are 15.77%, 15.89%, and 21.77% in Office, Baby, and Clothing, respectively. We credit this to the combined effects of the following three aspects. Firstly, we accurately extract structured semantic aspect-aware reviews using LLM to understand user behaviors at a finer level, which also avoids the noise in raw reviews. Secondly, the SAGCN performs graph convolution on various semantic-aspect-based graphs, which can alleviate the over-smoothing problem (Liu et al., 2021) and avoid negative information from unreliable neighbor nodes (Wei et al., 2020). Thirdly, we extract fine-grained user behaviors for each semantic aspect to bridge the modality gap between interaction and textual features.
4.3. Effect of Layer Numbers
(a) Recall on Office
(b) Recall on Clothing
Figure 4. Performance Comparison between SAGCN and competitors at different layers on Office and Clothing.
In this section, we evaluated the performance of our model by varying the number of layers in the graph convolution process. Our primary goal was to examine whether our model can mitigate the over-smoothing problem often encountered in graph convolutions. To investigate the effectiveness of our model, we iteratively increased the number of convolution layers from 1 to 7. We then compared its performance against four GCN-based methods: LightGCN, DGCF, IMP-GCN, GRCN, and BM3. The comparative outcomes are depicted in Figure 4.
From the results, all three models (DGCF, NCL, and BM3) achieve their peak performance when stacking 2 or 3 layers across two datasets. Beyond this, they will cause dramatic performance degradation with increasing layers. These results indicate that these three models suffer from the over-smoothing problem and inadvertently introduce noise information in the deep structure. GRCN achieves the peak performance when stacking 4 layers. This is attributed to discovering and pruning the potential false-positive edges, which avoid the noisy information from unreliable neighboring nodes. IMP-GCN can achieve robust performance with a deeper structure. This resilience can be credited to its effectiveness in filtering out noisy information from high-order neighbors. Our proposed SAGCN consistently outperforms all competitors when more than 2 layers are stacked, and it achieves peak performance when stacking 5 or 6 layers. This indicates that SAGCN can alleviate the over-smoothing problem by performing graph convolution over multiple semantic aspect-based graphs.
4.4. Semantic Aspects Analysis
4.4.1. Semantic Aspect-Aware Interactions Analysis
In our experiment, we only selected eight aspects per dataset due to sparse interactions in other aspects. Table 3 reported the number of interactions for different semantic aspects across three datasets.
Based on our observations of various recommendation scenarios, it is evident that users have distinct concerns regarding different aspects. For instance, in reviews of office products, convenience is a common concern, while fit is only mentioned in clothing reviews. To rank the semantic aspects in each recommendation scenario, we analyzed their interaction numbers. This helps us identify that the importance of different semantic aspects varies depending on the number of interactions. For example, in reviews of Office products, Durability was mentioned more frequently than Design. On the other hand, in reviews of Baby, Design was mentioned more frequently than Durability.
4.4.2. Effect of Aspect Numbers
(a) Recall on Office
(b) Recall on Clothing
Figure 5. Performance comparison of SAGCN with different semantic aspect numbers on Office and Clothing. Notice that the values are reported by percentage with ’%’ omitted.
To investigate the effect of semantic aspect numbers, we conducted experiments on Office and Clothing. Note that all aspects were ranked according to the number of interactions, we selected a number of aspects based on their ranking. Our findings, shown in Fig. 5, suggest that increasing the semantic aspect numbers can improve the performance of SAGCN. In other words, SAGCN achieves its peak performance when leveraging all aspects, indicating that it can benefit from various aspects. Furthermore, we found that SAGCN with eight aspects showed a greater improvement in performance than it with six aspects. This could be due to the fact that the added two aspects might provide more supplementary information to the other six. When combined, the full set of features gives the model a richer understanding of user behaviors.
4.5. Interpretability Analysis
(a) Recall on Office
(b) Recall on Clothing
Figure 6. Performance comparison of SAGCN on different semantic aspects on Office and Clothing. Notice that the values are reported by percentage with ’%’ omitted.
Figure 7. The preference score of user $u3547$ for item $i836$ and $i1322$ from different semantic aspects on Office. Notice that the values are reported by percentage with ’%’ omitted.
In this section, we examined the interpretability of our model from two different perspectives.
First, our model can interpret the contribution of each semantic aspect in a recommendation scenario. To examine the contribution of different aspects, we predicted the users’ preferences using individual aspect embeddings of the users and items. Fig. 6 shows the following interesting observations: (1) Different semantic aspect features have varying contributions in recommendation. For instance, Quality has a greater impact than Design for Office. (2) The number of semantic-aware interactions is not the main contribution factor. Surprisingly, even though the number of interactions for Price is the least, it still has the highest contribution to our model. This could be because Price introduces more valuable interactions than the other aspects.
Second, we can further provide a detailed explanation by providing the scores of various semantic aspects for each user-item pair. In Fig. 7, we computed the preference score of user $u3547$ for item $i836$ and $i1322$ from different semantic aspects on Office. From the result, we can conclude that the user would be interested in these two items mainly due to their Convenience and Quality. Additionally, the Comfort aspect of $i1322$ might also attract the user’s attention.
5. Conclusion
Our work proposes a novel approach that extracts semantic aspects from user reviews and utilizes them to enhance recommendation performance. Specifically, we devise a chain-based prompting approach with LLMs to extract the semantic aspect-aware reviews, which are then utilized to identify semantic aspect-aware user-item interactions. To improve the recommendation performance and interpretability, we propose a semantic aspect-based graph convolution network (SAGCN) that performs graph convolution operations on multiple semantic aspect-based graphs. We conducted extensive experiments on three benchmark datasets to evaluate the performance of our SAGCN model. Our experimental results demonstrate that SAGCN outperforms state-of-the-art GCN-based recommendation models. The experiment analysis exhibits the interpretability of our approach.
In summary, our work emphasizes the importance of analyzing user reviews by creating structured reviews to understand users before recommendations. On the other hand, we investigate a new approach to improving recommendation performance by utilizing LLMs.